# Visualize Geometric Series

### v2

@@ Abstract.
Key words and phrases:
Geometric series; proof without words; generalizations
2020 Mathematics Subject Classification:
40A05, 97-01
@@ 1. Introduction
Probably all calculus students have encountered the following formula to compute the geometric series $a+ar+ar^{2}+ar^{3}+\cdots$ when $|r|<1$:
@@
It is one of a few precious tools to find the exact sum of a convergent series. Other well-known tools include telescoping series and the Weierstrass factorization theorem, used by Euler in computing $\sum_{n=1}^{\infty}1/n^{2}$. It is worth emphasizing that (1.1) computes all geometric series as long as the common ratio has a magnitude less than $1$, but its usefulness does not stop there. For example, differentiating both sides of (1.1) with respect to $r$ gives another formula:
@@
While (1.1) is simply the Maclaurin series for $a(1-r)^{-1}$, a more elegant proof is to find a closed-form formula for the partial sums $s_{n}$ of the geometric series and then let $n$ go to infinity. In particular, observe that for $n\geq 1$,
@@
Though Mabry’s proof is only for the case $a=r=1/4$, it provides a nice intuition. If the area of the outermost triangle is $1$, then since the shaded triangles cover a third of each layer, their total area is $1/3$. Another way to compute the total shaded area is by summing up individual triangles: $1/4+(1/4)^{2}+(1/4)^{3}+\cdots$. The two ways must give the same answer, so we have (1.2).
@@
Edgar asked, “Is it possible to determine which other series allow analogous proof without words?”. The answer is, “It depends on what we mean by “analogous" proofs!". If the question asked for a series with $r=a$ as in Mabry’s and Edgar’s proofs, then we shall see that Figures 1 and 2 are unique of their kinds; if $r$ is allowed to be different from $a$, then there are many others.
We conclude this section with the beautiful proof without words by the Viewpoints 2000 Group [3]: Figure 3 proves (1.1) for all $r\in(0,1)$ rather than for special cases as Figures 1 and 2; however, we shall show that all of them share a common proof idea.
@@
Figure 3. Viewpoints 2000 Group’s proof: $a+ar+ar^{2}+\cdots=a/(1-r),\mbox{ for }0<r<1$.
@@ 2. Generalizing Mabry’s and Edgar’s proof
Figure 4 is the skeleton of Figures 1 and 2. The area of $\triangle{\textbf{ABC}}=1$, thus the length $|\textbf{BC}|=2\cdot 3^{-1/4}$. Let $n$ and $k$ be the number of triangles and of shaded triangles in each layer, respectively. Then $1\leq k<n$. Let $s\in(0,1)$ be the ratio of the lengths of consecutive parallel line segments. For example, $s$ is $1/2$ and $2/3$ in Figures 1 and 2, respectively.
$\scriptscriptstyle{\mathbf{B}}$$\scriptscriptstyle{\mathbf{C}}$$\scriptscriptstyle{2\cdot 3^{-1/4}}$$\scriptscriptstyle{\mathbf{A}}$$\scriptscriptstyle{2\cdot 3^{-1/4}s}$$\scriptscriptstyle{2\cdot 3^{-1/4}s^{2}}$$\scriptscriptstyle{2\cdot 3^{-1/4}s^{3}}$3rd layer2nd layer1st layer
@@
Another way to compute $\mathbf{T}$ is to sum up the areas of the shaded triangles. The area of the $1$st layer is $1-s^{2}$, meaning each triangle in the $1$st layer has area $u:=(1-s^{2})/n$. Due to similar triangles, each triangle in the $n$th layer has area $us^{2(n-1)}$ thanks to Property (P2). Therefore, the total shaded area is (recall that in each layer, $k$ triangles are shaded)
@@
As the two ways of computing $\mathbf{T}$ must produce the same result, we have the identity
@@
We are now ready to show that Mabry’s and Edgar’s proofs are the only ones that satisfy all the Properties (P1), (P2), and (P3). Indeed, their proofs require two conditions on $s$, which are simultaneously satisfied only when $s$ is $1/2$ or $2/3$.
@@
Furthermore, each triangle in the $1$st layer has area $(1-s)^{2}$, while the $1$st layer is $1-s^{2}$. Therefore, Property (P1) asks that $n=(1-s^{2})/(1-s)^{2}\in\mathbb{N}$, so $2/(1-s)\in\mathbb{N}$. Write $s=1-2/m$ for some integer $m\geq 3$. Then the number of triangles in each layer is $n=m-1$. Plugging this back to (2.2) gives
@@
The two restrictions on $s$ imply that $s\in\{1/2,2/3\}$. We conclude that Mabry’s and Edgar’s proofs are the only ones that satisfy all of the Properties (P1), (P2), and (P3).
If we drop Property (P3), then we have many other similar proofs to Figures 1 and 2. From the above analysis, we need only to choose $s=1-2/m$ for some $m\geq 3$ and partition each layer into $m-1$ equal triangles. However, this is only true when $m$ is even. Figure 5 considers the case $m=3$, a representative for odd $m$.
@@ 3. A common idea among the three proofs without words
At first glance, Figures 1, 2, and 3 are little related. However, they are based on the same idea. To illustrate this, we reposition the Viewpoints 2000 Group’s graph as in Figure 8.
@@
[…]

### v3

@@ 1. Introduction
Probably all calculus students have encountered the following formula to compute the sum of the geometric series $a+ar+ar^{2}+ar^{3}+\cdots$ when $|r|<1$:
@@
It is one of a few precious tools to find the exact sum of a convergent series. Other well-known tools include telescoping series and the Weierstrass factorization theorem, used by Euler in computing $\sum_{n=1}^{\infty}1/n^{2}$. It is worth emphasizing that (1.1) computes the sum of all geometric series as long as the common ratio has a magnitude less than 11, but its usefulness does not stop there. For example, differentiating both sides of (1.1) with respect to $r$ gives another formula:
@@
While (1.1) is simply the Maclaurin series for $a(1-r)^{-1}$, a more elegant proof is to find a closed form for the partial sums $s_{n}$ of the geometric series and then let $n$ go to infinity. In particular, observe that for $n\geq 1$,
@@
Though Mabry’s proof is only for the case $a=r=1/4$ in (1.1), it provides a nice intuition. If the area of the outermost triangle is 11, then since the shaded triangles cover a third of each layer, their total area is $1/3$. Another way to compute the total shaded area is by summing up individual triangles: $1/4+(1/4)^{2}+(1/4)^{3}+\cdots$. The two ways must give the same answer, so we have (1.2).
@@
There is a small difference between Figures 1 and 2. Figure 1 is divided into four triangles, then we add successive subdivisions to the upper $1/4$. Figure 2 is divided into nine triangles but we remove existing lines from the upper $4/9$ to divide that into nine triangles, then again leave only the bottom layer, and repeat the process.
Edgar asked, “Is it possible to determine which other series allow analogous proof without words?”. The answer is, “It depends on what we mean by “analogous” proofs!”. If the question asked for a series with $r=a$ as in Mabry’s and Edgar’s proofs, then we shall see that Figures 1 and 2 are unique of their kinds; if $r$ is allowed to be different from $a$, then there are many others.
We conclude this section with the beautiful proof without words by the Viewpoints 2000 Group [3]: Figure 1.4 proves (1.1) for all $r\in(0,1)$ rather than for special cases as Figures 1 and 2, and $a$ and $r$ are not necessarily equal; however, we shall show that all of them share a common proof idea.
@@
Figure 3. Viewpoints 2000 Group’s proof:
$a+ar+ar^{2}+\cdots\ =\ \frac{a}{1-r},\mbox{ for }0<r<1.$ (1.4)
The clever idea of Figure 1.4 lies in the slopes of the two lines $y=x$ and $y=rx+a$ so that moving horizontally preserves the previous distance, while moving vertically shrinks the previous distance by a factor of $r$. Hence, the sum of vertical distances produces the wanted series, $a+ar+ar^{2}+\cdots$.
@@ 2. Generalizing Mabry’s and Edgar’s proof
Figure 4 is the skeleton of Figures 1 and 2. The area of $\triangle{\textbf{ABC}}=1$, thus the length $|\textbf{BC}|=2\cdot 3^{-1/4}$. (Can you use the geometry of an equilateral triangle to see why this is true?) Let $n$ and $k$ be the number of triangles and of shaded triangles in each layer, respectively. Then $1\leq k<n$. Let $s\in(0,1)$ be the ratio of the lengths of consecutive parallel line segments. For example, $s$ is $1/2$ and $2/3$ in Figures 1 and 2, respectively.
$\scriptscriptstyle{\mathbf{B}}$$\scriptscriptstyle{\mathbf{C}}$$\scriptscriptstyle{2\cdot 3^{-1/4}}$$\scriptscriptstyle{\mathbf{A}}$$\scriptscriptstyle{2\cdot 3^{-1/4}s}$$\scriptscriptstyle{\mathbf{D}}$$\scriptscriptstyle{\mathbf{E}}$$\scriptscriptstyle{2\cdot 3^{-1/4}s^{2}}$$\scriptscriptstyle{2\cdot 3^{-1/4}s^{3}}$3rd layer2nd layer1st layer
@@
Another way to compute $\mathbf{T}$ is to sum up the areas of the shaded triangles. Since $|\mathbf{DE}|=s|\mathbf{BC}|$, the area of $\triangle\textbf{ADE}=s^{2}\triangle\textbf{ABC}=s^{2}$. Hence, the area of the 11st layer is $1-s^{2}$, meaning each triangle in the 11st layer has area $u:=(1-s^{2})/n$.
Property (P2) guarantees that the edge ratio of two triangles from adjacent layers is $s$. Thus, moving from one layer to the immediate layer above shrinks the triangle area by $s^{2}$. It follows that each triangle in the $j$th layer has area $us^{2(j-1)}$. Therefore, the total shaded area is (recall that in each layer, $k$ triangles are shaded)
@@
As the two ways of computing $\mathbf{T}$ must produce the same result, we have the following identity.
Lemma 2.1.
For integers $1\leq k<n$ and $s\in(0,1)$. If $u=(1-s^{2})/n$, then
@@
We are now ready to prove our main result.
Theorem $2.$
Mabry’s and Edgar’s proofs are the only ones that satisfy all the Properties (P1), (P2), and (P3).
Proof.
We shall show that Mabry’s and Edgar’s proofs require two conditions on $s$, which are simultaneously satisfied only when $s$ is $1/2$ or $2/3$.
@@
That $|\textbf{DE}|=s|\textbf{BC}|$ implies that the height of a triangle in the first layer is (1s)(1-s) times the height of $\triangle\textbf{ABC}$. Hence, each triangle in the 11st layer has area $(1-s)^{2}$, while the 11st layer has area $1-s^{2}$. Therefore, Property (P1) asks that $n=(1-s^{2})/(1-s)^{2}\in\mathbb{N}$, so
$\frac{2}{1-s}\ =\ \frac{1-s^{2}}{(1-s)^{2}}+1\in\mathbb{N}.$
Write s=12/ms=1-2/m for some integer $m\geq 3$. Then the number of triangles in each layer is n=m1n=m-1. Plugging this back to (2.2) gives
@@
The two restrictions on $s$ imply that $s\in\{1/2,2/3\}$. We conclude that Mabry’s and Edgar’s proofs are the only ones that satisfy all of the Properties (P1), (P2), and (P3). ∎
If we drop Property (P3), then we have many other similar proofs to Figures 1 and 2. From the above analysis, we need only to choose s=12/ms=1-2/m for some $m\geq 3$ and partition each layer into m1m-1 equal triangles. However, this is only true when $m$ is even. Figure 5 considers the case $m=3$, a representative for odd $m$.
@@
Unlike Mabry’s and Edgar’s proofs, Figures 6 and 7 prove particular cases of Equation (2.1) when $ku\neq s^{2}$.
@@ 3. A common idea among the three proofs without words
At first glance, Figures 1, 2, and 1.4 are little related. However, they are based on the same idea. To illustrate this, we reposition the Viewpoints 2000 Group’s graph as in Figure 8. We shall prove (1.1) for $a=1$ and $r\in(0,1)$. This allows us to start with a triangle of base 11 instead of $a$ as in Figure 1.4.
@@
On the other hand, the shaded area covers $(1+\sqrt{r})^{-1}$ of the area of each layer. Hence, the shaded area covers $(1+\sqrt{r})^{-1}$ of the area of $\triangle{\textbf{ABC}}$. Since $\triangle\textbf{ABC}$ has area $\frac{1}{2(1-\sqrt{r})}$, we have
@@
Acknowledgement
The authors would like to thank the anonymous referees for various helpful suggestions.